Question
A researcher wants a 72 ounce saline solution with a concentration of 7%. She has concentrations of 2% and 14%. How much of each solution must she use to get the desired concentration?

Answer by ikleyn(52786)

.
The equation is

0.02*x + 0.14*(72-x) = 0.07*72.   (x = the volume of the 2% solution, in mL).



You need to understand why it is so, but is is SO EASY . . . !

    Both sides of the equation represent the balance of the salt mass in the solutions.



To solve the equation,  multiply its both sides by 100 to get it in more easy form. You will get

2x + 14(72 -x) = 504.


Now simplify and solve for x.


I leave it to you to complete the solution.


There is entire bunch of introductory lessons covering various types of mixture problems
    - Mixture problems
    - More Mixture problems
    - Solving typical word problems on mixtures for solutions
    - Word problems on mixtures for antifreeze solutions
    - Word problems on mixtures for alloys
    - Typical word problems on mixtures from the archive
in this site.

Read them and become an expert in solution the mixture word problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook in the section "Word problems" under the topic "Mixture problems".